# Kennedy-Tasaki transformation and non-invertible symmetry in lattice models beyond one dimension

### v1

@@ D.1 One dimension
[…]
@@
$\text{KT}_{l,N}Z_{2i-1}^{l}X_{2i}(Z_{2i+1}^{\dagger})^{l}$ $=Z_{2i}^{\dagger}Z_{2i+2}\text{KT}_{l,N},$ (185a)
$\text{KT}_{l,N}(Z_{2i}^{\dagger})^{l}X_{2i+1}Z_{2i+2}^{l}$ $=Z_{2i+1}^{\dagger}Z_{2i+3}\text{KT}_{l,N}.$ (185b)
@@ D.2 Two and higher dimensions
In our case $G_{s}=\mathbb{Z}_{N}\times\mathbb{Z}_{N}$ and the SSPT classification give $\mathcal{C}[\mathbb{Z}_{N}\times\mathbb{Z}_{N}]=\mathbb{Z}_{N}\times\mathbb{Z}_{N}\times\mathbb{Z}_{N}$. To enumerate various SSPT phases, we consider two square lattices with color red and blue as in Fig. 3(a). The generators for the three $\mathbb{Z}_{N}$ factors are 1) $\mathbb{Z}_{N}\times\mathbb{Z}_{N}$ SSPT with cluster entangler between adjacent red and blue sites, 2) $\mathbb{Z}_{N}$ SSPT (generalization of $\mathbb{Z}_{2}$ SSPT we considered in two dimensions in section VI) on red lattice, and 3) $\mathbb{Z}_{N}$ SSPT on blue lattice. It should be straightforward to generalize the KT transformation to each of the three generators of SSPT phase resulting in 1) two copies of plaquette clock model, 2) one copy of double-plaquette clock model, and 3) one copy of double-plaquette clock model, respectively.
@@ Appendix E The $\mathbb{Z}_{2}$ SSPT
On a 2d square lattice, where each site contains 2 qubits, one can define the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SSPT for subsystem symmetries. The SSPT state is equivalent to the 2d cluster state given by Hamiltonian
@@
After inserting the symmetry breaking terms, the Hamiltonian becomes
@@
In particular when $\lambda=0$, we directly show here that the $\beta(g)$ phase defined in Ref. [9] is nontrivial. For an SSPT state, because of its short-ranged entangled nature, the truncated symmetry operator $U^{j_{0},j_{1}}_{i_{0},i_{1}}\equiv\prod_{i=i_{0},j=j_{0}}^{i_{1},j_{1}}X$ only affects locally on the corners,
@@
where $S_{i}\equiv\prod_{i^{\prime}\geq i}\prod_{j}X_{i,j}$ is the symmetry action on a half-plane. Therefore, the ground state when $\lambda=0$ is a strong SSPT protected by a linearly-symmetric local unitary (LSLU). We note that a similar calculation can be performed for the 3D $\mathbb{Z}_{2}$ SSPT model.

### v2

@@ D.1 One dimension
$\begin{split}H^{l,N}_{1DSPT}&=-\sum_{i=1}^{L}\left(\left(Z_{2i-2}^{\dagger}\right)^{l}X_{2i-1}Z_{2i}^{l}+Z_{2i-1}^{l}X_{2i}\left(Z_{2i+1}^{\dagger}\right)^{l}+\text{h.c}\right).\end{split}$ (173)
@@
$\text{KT}_{l,N}Z_{2i-1}^{l}X_{2i}(Z_{2i+1}^{\dagger})^{l}$ $=Z_{2i}Z_{2i+2}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}\text{KT}_{l,N},$ (178a)
$\text{KT}_{l,N}(Z_{2i}^{\dagger})^{l}X_{2i+1}Z_{2i+2}^{l}$ $=Z_{2i+1}Z_{2i+3}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}\text{KT}_{l,N}.$ (178b)
@@ D.2 Two and higher dimensions
In our case $G_{s}=\mathbb{Z}_{N}\times\mathbb{Z}_{N}$ and the SSPT classification give $\mathcal{C}[\mathbb{Z}_{N}\times\mathbb{Z}_{N}]=\mathbb{Z}_{N}\times\mathbb{Z}_{N}\times\mathbb{Z}_{N}$. To enumerate various SSPT phases, we consider two square lattices with the color red and blue as in Fig. 4(a). The generators for the three $\mathbb{Z}_{N}$ factors are 1) $\mathbb{Z}_{N}\times\mathbb{Z}_{N}$ SSPT with cluster entangler between adjacent red and blue sites, 2) $\mathbb{Z}_{N}$ SSPT (generalization of $\mathbb{Z}_{2}$ SSPT we considered in two dimensions in section VI) on red lattice, and 3) $\mathbb{Z}_{N}$ SSPT on blue lattice. It should be straightforward to generalize the KT transformation to each of the three generators of the SSPT phase resulting in 1) two copies of the plaquette clock model, 2) one copy of the double-plaquette clock model, and 3) one copy of the double-plaquette clock model, respectively.
@@ Appendix E The $\mathbb{Z}_{2}$ SSPT
On a 2d square lattice, where each site contains 2 qubits, one can define the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SSPT for subsystem symmetries. The SSPT state is equivalent to the 2d cluster state given by the Hamiltonian
@@
After inserting the symmetry-breaking terms, the Hamiltonian becomes
@@
In particular, when $\lambda=0$, we directly show here that the $\beta(g)$ phase defined in Ref. [9] is nontrivial. For an SSPT state, because of its short-ranged entangled nature, the truncated symmetry operator $U^{j_{0},j_{1}}_{i_{0},i_{1}}\equiv\prod_{i=i_{0},j=j_{0}}^{i_{1},j_{1}}X_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,j}}$ only affects locally on the corners,
@@
where $S_{i}\equiv\prod_{i^{\prime}\geq i}\prod_{j}X_{i^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}^{\prime}}},j}$ is the symmetry action on a half-plane. Therefore, the ground state when $\lambda=0$ is a strong SSPT protected by a linearly symmetric local unitary (LSLU). We note that a similar calculation can be performed for the 3D $\mathbb{Z}_{2}$ SSPT model.
Appendix F Measurement-based gauging method
In this section, we provide a measurement-based method to gauge the symmetries of various models that we consider in the main text in sections III and VI.
F.1 Plaquette Ising model
The plaquette Ising model has both horizontal and vertical subsystem symmetries. To gauge these subsystem symmetries, we introduce an ancilla d.o.f. in $\ket{+}$ state at each center of a plaquette $p\in\Delta_{p}$ (a vertex of the dual lattice). This d.o.f. is entangled to the vertices ($v\in\Delta_{v}$) of the corresponding plaquette via controlled-Z gate as shown in Fig. 9. Then, we measure away the vertex d.o.f. in the Pauli X basis with the measurement outcome being all $\ket{+}$ state. Explicitly, this procedure can be written as an operator
$\text{KW}=\bra{+}^{\otimes\Delta_{v}}\prod_{p}\prod_{v\in\partial p}CZ_{p,v}\ket{+}^{\otimes\Delta_{p}}.$ (188)
Then, KW satisfies
$X_{v}=\prod_{p,v\in\partial p}Z_{p}\text{KW},$ KW , (189)
$\prod_{v\in\partial p}Z_{v}=X_{p}\text{KW},$ KW , (190)
mapping between the plaquette Ising models on the original and dual lattices. This allows a short-depth implementation of the KW transformation and can be easily generalized to higher dimensions. When measurement outcomes do not appear as the product state |+ket|$\rangle$| + ⟩, one can perform feedforward corrections to clean up non-trivial effects on the output wave function caused by them. Such a mechanism was explained in Refs. [49, 59] for gauging 0-form symmetries, and will be explained in Ref. [60] for more general spin models.
Figure 9: Cluster entanglement pattern for the KW duality map that maps between two plaquette Ising models. The blue lattice is the original lattice, and the red lattice is the ancilla lattice.
F.2 Double plaquette Ising model
In addition to the horizontal and vertical symmetries, the double plaquette Ising model also has a diagonal symmetry. To implement the self-dual transformation of the double plaquette Ising model, we introduce another copy of the square lattice on top of the original lattice which we call the ancilla lattice. Let us denote the superimposed vertices on the two lattices by $v$ and $v^{(a)}$, where the superscript $(a)$ denotes the lattice on which the ancilla d.o.f. is initialized in the $\ket{+}$ state. A particular $v^{(a)}$ is entangled via controlled-Z gate to adjacent vertices on the original lattice as shown in Fig. 10. Let us denote the set of neighboring vertices (including diagonal but not off-diagonal) of an ancilla vertex $v^{(a)}$ by $\text{Neigh}(v^{(a)})$ and the set of neighboring ancilla vertices (including diagonal but not off-diagonal) of a vertex $v$ by $\text{Neigh}(v)$. Then, we measure away the vertex d.o.f. in the original lattice in the Pauli X basis with measurement outcome being all $\ket{+}$ state. Explicitly, this procedure is given by
$\text{KW}=\bra{+}^{\otimes\Delta_{v}}\prod_{v^{(a)}}\prod_{v\in\text{Neigh}(v^{(a)})}CZ_{v^{(a)},v}\ket{+}^{\otimes\Delta_{v^{(a)}}}$ (191)
KW satisfies
$X_{v}=\prod_{v^{(a)}\in\text{Neigh}(v)}Z_{v^{(a)}}\text{KW},$ KW , (192)
$\prod_{v\in\text{Neigh}(v^{(a)})}Z_{v}=X_{v^{(a)}}\text{KW}.$ KW . (193)
mapping between the double plaquette Ising models on the original and ancilla lattices. This allows us to perform the KW duality and can be generalized to higher dimensions. Similar to the discussion in the plaquette Ising model, when measurement outcomes do not appear as the product state $\ket{+}$, one can perform the feedforward corrections to clean up the non-trivial effect on the output wavefunction caused by them [60].
Figure 10: Cluster entanglement pattern for the KW duality map that maps between two double-plaquette Ising models. The blue lattice is the original lattice, and the red lattice is the ancilla lattice.
F.3 Implementing the KT transformation
We have seen in the previous subsections how to implement the KW transformations. Equipped with this, we can then use the duality web to perform the KT transformation as well by composing, e.g., $TST$, where $T$ can be implemented by the cluster-state entangling operation (and $S$ is the KW operation).
Appendix G Mapping of order parameters
Order parameters of non-trivial phases get mapped under the KT transformations. Let $B_{k}$ be a product of $Z$ operators in the models we have considered in this work exhibiting SSSB phases, such as the plaquette Ising term, the double plaquette Ising term, etc. Let $K_{k}$ be the stabilizer of the corresponding SSPT cluster states. In SSSB phases, as a generalization of the two-point function $\langle Z_{j}Z_{j+\ell}\rangle$ in the 1d Ising model, a long-range entanglement can be detected by the order parameter
$\langle\prod_{k\in\mathcal{M}}B_{k}\rangle,$ (194)
where $\mathcal{M}$ is a segment consisting of cells (hypercubes) along a line where the subsystem symmetries are supported. An example in the 2d plaquette Ising model is the product of $B_{k}=Z_{i,j}Z_{i+1,j}Z_{i+1,j}Z_{i+1,j+1}$ along a 1d segment (a thin strip consisting of plaquettes), which becomes a product of four Pauli $Z$’s at the corners of the strip: $\langle Z_{i,j}Z_{i+\ell,j}Z_{i+\ell,j+1}Z_{i,j+1}\rangle$. Upon the KT transformation, the corresponding SSPT string order parameter is written as
$\langle\prod_{k\in\mathcal{M}}K_{k}\rangle,$ (195)
where the segment $\mathcal{M}$ is filled with $X$ operators in addition to the $Z$ operators at the corners. For the 2d plaquette Ising model, an example of the string order parameter obtained by the KT operator (50) is $\langle Z_{i,j}Z_{i,j+1}X_{i+\frac{1}{2},j+\frac{1}{2}}\cdots X_{i+\ell-\frac{1}{2},j+\frac{1}{2}}Z_{i+\ell,j}Z_{i+\ell,j+1}\rangle$.